Give an example of infinite ring of characteristic 7.
1
Expert's answer
2020-10-20T17:52:13-0400

Consider the ring "\Bbb Z_7[x]" of polynomials in one variable 𝑥 with coefficients in ℤ"_7" It is an infinite ring since "x^m \in \Bbb Z_7[x]:=\{a_1x^m+a_2x^{m-1}+...+a_n:a_1\ne 0,a_1,...,a_n\in \Bbb Z_7\& m\in \Bbb Z^+\}" and note that for all positive integers 𝑚, and "x^{m_1} \ne x^{m_2}" for "m_1\ne m_2" ,Thus "Cardinality( \Bbb Z_7[x])=\infty" But the charactetistic of "\Bbb Z_7[x]" is 7 as 7 is prime and 1+1+1+1+1+1+1=7=0. in the ring "\Bbb Z_7" .